OpenStudy (anonymous):

Find lim x approaches 0 of sin(2x)/x using only sanwich thorem or limit properties. Thnaks

OpenStudy (zarkon):

\[\frac{\sin(2x)}{x}=2\frac{\sin(2x)}{2x}\]

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

can go further

OpenStudy (zarkon):

\[\lim_{x\to0}\frac{\sin(2x)}{x}=\lim_{x\to0}2\frac{\sin(2x)}{2x}=2\cdot 1=2\]

OpenStudy (anonymous):

how do we know that sin(2x)/2x is 1?

OpenStudy (angela210793):

its called ''important limit''